Line 'L' has interceps a and b on the coordinate axis. When the axis are rotated through a given angle, keeping the origin fixed, the same line 'L' has intercepts p and q on the new axis, then


A) a2 + b2 = p2 + q2


B) a2 + p2 = b2 + q2


C) 1/a2 + 1/b2 = 1/p2 + 1/q2


D) 1/a2 + 1/p2 = 1/b2 + 1/q2

Grade:12

4 Answers

Pratham Ashish
14 years ago

hi,

 in the first coordinate sys,

x/a + y/b = 1     ..............................1

in other

x'/ p + y'/q  =1

where

 

\begin{align} x'&=x\cos\theta+y\sin\theta\\ y'&=-x\sin\theta+y\cos\theta. \end{align}

 

so,  x {  cos θ/p - sinθ /q }   + y { sinθ /p  + cosθ/ q }  =1 

comparing with (1).....

 {  cos θ/p - sinθ /q }/ 1/a   = { sinθ /p  + cosθ/ q }  / 1/b = 1/1

 cos θ/p - sinθ /q   = 1/a  ...............2

sinθ /p  + cosθ/ q  = 1/b ........................3

sq. and adding 2 &3

1/p^2 +  1/q^2  = 1/a^2 + 1/b^2

 

 

 

 

 

 

 

Aditya Lather
5 years ago
H
Aditya Lather
5 years ago
Here the length of perpendicular from origin to the lines would be equal hence on equating that distance we get
√(1/p^2 +1/q^2)=√(1/a^2+1/b^2)
On squaring both sides we get the answer
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
Here the length of perpendicular from origin to the lines would be equal hence on equating that distance we get
√(1/p^2 +1/q^2)=√(1/a^2+1/b^2)
On squaring both sides we get the answer
Hence the correct option will be C
 
Thanks and regards,
Kushagra